### v1

# Strassen $2\times 2$ Matrix Multiplication from a 3-dimensional Volume Form

Abstract
The Strassen $2\times 2$ matrix multiplication algorithm arises from the volume form on the 3-dimensional quotient space of the $2\times 2$ matrices by the multiples of identity.
1 Introduction
Strassen’s $2\times 2$ matrix multiplication algorithm [1] is a formula for multiplying $2\times 2$ matrices $a$ and $b$:
$ab=\mathrm{tr}(a)\mathrm{tr}(b)I+\sum_{i=1}^{6}\mathrm{tr}(X_{i}a)\mathrm{tr}(Y_{i}b)Z_{i}$ (1)
where $I$ is the identity matrix, $\mathrm{tr}$ is the trace, and the $X_{i},Y_{i},Z_{i}$ are constant matrices. This formula is a rank 7 decomposition of the matrix multiplication tensor, that is, a decomposition of matrix multiplication into a sum of 7 simple tensors.
This may be applied recursively to multiply $n\times n$ matrices in $O(n^{\log 7/\log 2})$ time, approximately $O(n^{2.81})$, opening a research field to which the book [2] provides an introduction. One line of research has focused on further improving this asymptotic complexity, notably [3], [4] achieving $O(n^{2.376})$ and several refinements, recently [5] and [6], approaching $O(n^{2.37})$. Another line of research has pursued decompositions of matrix multiplication tensors for other small matrix sizes such as $3\times 3$, $4\times 4$, etc. These have often involved numerical searches, such as the recent [7].
Despite these advances, matrix multiplication algorithms faster than $O(n^{2.81})$ are “almost never implemented” [8], and practical evaluations such as [9] have continued favoring the Strassen $2\times 2$ algorithm. Known algorithms for other small matrix sizes struggle to significantly improve on it: the up-to-date table [10] shows complexity exponents clustering around 2.8. For $4\times 4$ matrix multiplication, the Strassen algorithm has tensor rank $497^{2}=$ 49, and that remained the state of the art for over 55 years until [7] and [11] lowered that from 49 to 48, achieving a complexity exponent of $2.79$. Moreover, known algorithms with substantially lower asymptotic complexity tend to have large constants in the $O$, as discussed in [12].
The Strassen $2\times 2$ algorithm also stands out from a theoretical perspective: its tensor rank 7 is known to be optimal [13] and it is known to be essentially unique under that constraint [14]. By contrast, the tensor rank of $n\times n$ matrix multiplication is still unknown for all $n\geq 3$. For $n=3$, it is still only known to be between 19 and 23, see [15].
Optimality and uniqueness make the Strassen $2\times 2$ algorithm a basic fact of 2-dimensional linear algebra. Such facts are expected to be simple and geometric. However, the original statement and proof of the Strassen algorithm are calculations on matrix coefficients. This has motivated a quest for geometric interpretations. The recent [16] and [8] in particular were inspirational to the present article, and [8] contains a survey of this endeavour, tracing it back to the years following the publication of the original Strassen article [1]. Other recent articles in this line of research include [17], [18], [19] and [20].
The present article offers a geometric interpretation of the Strassen algorithm by addressing a more general question: is the Strassen algorithm an independent fact in multilinear algebra, or could it be related to a known fact? We derive it from the expansion of a 3-dimensional volume form into an antisymmetrized sum of $3!=6$ simple tensors. That expansion follows from the one-dimensionality of the space of antisymmetric $n$-forms, which is an abstract version of Cavalieri’s principle, the idea that the volume of a solid is unchanged by sliding parallel slices. As to the question of why specifically $2\times 2$ matrices, the answer is that as matrix multiplication is a tensor of order 3 on matrix spaces, interpreting it as a volume form requires a 3-dimensional matrix space, and the specific case of $2\times 2$ matrices gives us such a 3-dimensional matrix space by taking the quotient by multiples of the identity matrix: $13=2^{2}-$ 1.
Acknowledgements. The author would like to thank Paolo d’Alberto and Zach Garvey at AMD for helpful comments.
2 Overview
Fix, for this entire article, a 2-dimensional vector space $V$ over a field $k$. Let $\mathrm{L}(V)$ denote the space of linear maps from $V$ to itself. Start by considering this trilinear form $g$ on $\mathrm{L}(V)$:
$g(a_{1},a_{2},a_{3})=\mathrm{tr}(a_{1}a_{2}a_{3})-\mathrm{tr}(a_{3}a_{2}a_{1}).$ (2)
We notice (Lemma 12) that $g$ is a volume form on the quotient of $\mathrm{L}(V)$ by the multiples of the identity matrix, which has dimension 3. This gives (Lemma 13) rank 6 decompositions of $g$ parametrized by bases of the dual space. Our next step is to relate $g$ to this other trilinear form $h$ on $\mathrm{L}(V)$:
$h(a_{1},a_{2},a_{3})=\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})-\mathrm{tr}(a_{1}a_{2}a_{3}).$ (3)
Using the natural isomorphism $\mathrm{L}(V)^{*\otimes 3}\simeq\mathrm{L}(V^{\otimes 3})$, view the trilinear forms $\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})$, $\mathrm{tr}(a_{1}a_{2}a_{3})$ and $\mathrm{tr}(a_{3}a_{2}a_{1})$ as respectively the permutations $\mathrm{id}$, (123) and (321) permuting the terms in $V^{\otimes 3}$ (Lemma 8). This allows viewing $h$ as the composition of $g$ with a linear map induced by the permutation (321) (Lemma 16), which allows transporting certain rank 6 decompositions of $g$ into rank 6 decompositions of $h$ (Proposition 18, our main result), yielding (Corollary 20)
$\mathrm{tr}(a_{1}a_{2}a_{3})=\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})-\textrm{\{rank 6 decomposition of $h$\}},$ (4)
which is a rank 7 decomposition of $\mathrm{tr}(a_{1}a_{2}a_{3})$. Dualizing that yields a rank 7 decomposition of matrix multiplication (Corollary 21) parametrized by a choice of basis. A specific choice yields the original Strassen algorithm (Corollary 22).
3 Terminology and lemmas in tensor algebra
Throughout this article, “vector space” means finite-dimensional vector space. For any vector spaces $U$ and $W$ over a field $k$, let $\mathrm{L}(U,W)$ denote the space of linear maps from $U$ to $W$. In the case $W=U$, we write $\mathrm{L}(U)$ for $\mathrm{L}(U,U)$. In the case $W=k$, we let $U^{*}=\mathrm{L}(U,k)$ denote the dual space of $U$.
Given any vector spaces $U_{1},\ldots,U_{n}$, $W_{1},\ldots,W_{n}$, we will make the identification
$\mathrm{L}(U_{1},W_{1})\otimes...\otimes\mathrm{L}(U_{n},W_{n})\simeq\mathrm{L}(U_{1}\otimes...\otimes U_{n},W_{1}\otimes...\otimes W_{n}).$
As special cases of that, for any vector space $U$ over $k$, for any positive integer $n$, we identify $\mathrm{L}(U)^{\otimes n}\simeq\mathrm{L}(U^{\otimes n})$ and $U^{*\otimes n}\simeq(U^{\otimes n})^{*}$. The latter identification means concretely that given linear forms $\mu_{1},\ldots,\mu_{n}$ on a vector space $U$, we identify the tensor $\mu_{1}\otimes\ldots\otimes\mu_{n}$ as the $n$-linear form on $U$ given, for all vectors $u_{1},\ldots,u_{n}$ in $U$, by:
$(\mu_{1}\otimes\ldots\otimes\mu_{n})(u_{1},\ldots,u_{n})=\mu_{1}(u_{1})\ldots\mu_{n}(u_{n}).$
Let us now describe a few other natural isomorphisms of tensor spaces that we will keep as named isomorphisms, refraining from making identifications.
Definition 1.
For any vector space $U$, define linear maps $\iota$, $\iota^{*}$ and *:
$\iota\;$ $:\;U\otimes U^{*}\rightarrow\mathrm{L}(U),$ $v\otimes\lambda$ $\mapsto\iota(v\otimes\lambda)=(u\mapsto\lambda(u)v)$
$\iota^{*}\;$ $:\;U\otimes U^{*}\rightarrow\mathrm{L}(U)^{*},$ $v\otimes\lambda$ $\mapsto\iota^{*}(v\otimes\lambda)=(a\mapsto\lambda(a(v)))$
*\; $:\;\mathrm{L}(U)\rightarrow\mathrm{L}(U)^{*},$ $a$ $\mapsto a^{*}=(b\mapsto\mathrm{tr}(ab)).$
Lemma 2.
The linear maps $\iota$, $\iota^{*}$ and * are isomorphisms.
Proof.
These maps are injective, and when $U$ has dimension $n$, the source and destination spaces have the same dimension $n^{2}$. ∎
Lemma 3.
For any vector space $U$, for any $u,v$ in $U$ and any $\lambda,\mu$ in $U^{*}$, we have
$\iota(v\otimes\lambda)\iota(u\otimes\mu)$ $=\lambda(u)\iota(v\otimes\mu),$ (5)
$\iota^{*}(v\otimes\lambda)(\iota(u\otimes\mu))$ $=\lambda(u)\mu(v),$ (6)
$\mathrm{tr}(\iota(v\otimes\lambda))$ $=\lambda(v).$ (7)
Proof.
For any $w$ in $U$, we have $\iota(v\otimes\lambda)\iota(u\otimes\mu)(w)=\iota(v\otimes\lambda)(\mu(w)u)=\lambda(u)\mu(w)v=\lambda(u)\iota(v\otimes\mu)(w)$, establishing Equation (5). We have $\iota^{*}(v\otimes\lambda)(\iota(u\otimes\mu))=\lambda(\iota(u\otimes\mu)(v))=\lambda(\mu(v)u)=\mu(v)\lambda(u)$, establishing Equation (6). Let $w_{1},\ldots,w_{n}$ be a basis of $U$ such that $w_{1}=v$. In that basis, the matrix of $\iota(v\otimes\lambda)$ is $\left(\begin{smallmatrix}\lambda(v)&\lambda(w_{2})&\cdots&\lambda(w_{n})\\ 0&0&\cdots&0\\ \cdots&\cdots&\cdots&\cdots\end{smallmatrix}\right),$ whose trace is $\lambda(v)$, establishing Equation (7). ∎
Lemma 4.
For any vector space $U$, the following diagram commutes, justifying the notation $\iota^{*}$.
[…]
Proof.
The claim is that for all $v\otimes\lambda$ in $U\otimes U^{*}$, we have $\iota^{*}(v\otimes\lambda)=\iota(v\otimes\lambda)^{*}$ as elements of $\mathrm{L}(U)^{*}$. By linearity, it is enough to check that these forms in $\mathrm{L}(U)^{*}$ agree on rank one elements of $\mathrm{L}(U)$, which are the $\iota(u\otimes\mu)$ with $u$ in $U$ and $\mu$ in $U^{*}$. Indeed, the equations from Lemma 3 give:
$\iota^{*}(v\otimes\lambda)(\iota(u\otimes\mu))$ $=\lambda(u)\mu(v)$ by Equation (6)
$=\mathrm{tr}(\lambda(u)\iota(v\otimes\mu))$ by Equation (7)
$=\mathrm{tr}(\iota(v\otimes\lambda)\iota(u\otimes\mu))$ by Equation (5)
$=\iota(v\otimes\lambda)^{*}(\iota(u\otimes\mu))$ by Definition 1. $\;\qed$
Definition 5.
For any vector space $U$ and any element $a$ of $\mathrm{L}(U)$, let $L_{a},R_{a}$ denote respectively the left and right multiplication-by-$a$ maps: $L_{a}(b)=ab$ and $R_{a}(b)=ba$ for all $b$ in $\mathrm{L}(U)$. Thus $L_{a}$ and $R_{a}$ are elements of $\mathrm{L}(\mathrm{L}(U))$.
Lemma 6.
For any vector space $U$ over a field $k$ and any $a,b$ in $\mathrm{L}(U)$, we have
$(ab)^{*}=a^{*}\circ L_{b}=b^{*}\circ R_{a}$
where $\circ$ denotes the composition of linear maps $\mathrm{L}(U)\rightarrow\mathrm{L}(U)\rightarrow k$.
Proof.
For any $c$ in $\mathrm{L}(U)$, we have
$(ab)^{*}(c)=\mathrm{tr}(abc)=\mathrm{tr}(aL_{b}(c))=a^{*}(L_{b}(c))=(a^{*}\circ L_{b})(c)$
and similarly
$(ab)^{*}(c)=\mathrm{tr}(abc)=\mathrm{tr}(bca)=\mathrm{tr}(bR_{a}(c))=b^{*}(R_{a}(c))=(b^{*}\circ R_{a})(c).\qed$
Definition 7.
For any vector space $U$, for any permutation $\sigma$ in the symmetric group $S_{3}$, define a map $t_{\sigma}$ in $\mathrm{L}(U^{\otimes 3})$ by letting, for all $u_{1},u_{2},u_{3}$ in $U$,
$t_{\sigma}(u_{1}\otimes u_{2}\otimes u_{3})=u_{\sigma(1)}\otimes u_{\sigma(2)}\otimes u_{\sigma(3)}.$ (9)
The following lemma is classical and is encountered around Weyl invariant tensor theory ([21], [22]), which, among other things, establishes that the $t_{\sigma}$ span the space of $\mathrm{GL}(U)$-invariant tensors in $\mathrm{L}(U^{\otimes 3})$. We will not need any of that theory, but it is still useful context.
Lemma 8.
For any vector space $U$, the images of $t_{\mathrm{id}}$, $t_{(123)}$, $t_{(321)}$ under the map $t\mapsto t^{*}$ from Definition 1 are the following trilinear forms, given by their values at any $a_{1}\otimes a_{2}\otimes a_{3}$ in $\mathrm{L}(U)^{\otimes 3}$:
$t_{\mathrm{id}}^{*}(a_{1}\otimes a_{2}\otimes a_{3})$ $=$ $\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3}),$
$t_{(123)}^{*}(a_{1}\otimes a_{2}\otimes a_{3})$ $=$ $\mathrm{tr}(a_{1}a_{2}a_{3}),$
$t_{(321)}^{*}(a_{1}\otimes a_{2}\otimes a_{3})$ $=$ $\mathrm{tr}(a_{3}a_{2}a_{1}).$
Proof.
By linearity, it is enough to prove these equations in the case where the $a_{i}$ are simple tensors of the form $a_{i}=\iota(v_{i}\otimes\lambda_{i})$ with $v_{i}$ in $U$ and $\lambda_{i}$ in $U^{*}$. Letting the dot ($\cdot$) denote multiplication in $\mathrm{L}(U^{\otimes 3})$, for any permutation $\sigma$ in $S_{3}$, we have
$t_{\sigma}^{*}(a_{1}\otimes a_{2}\otimes a_{3})$ $=$ $\mathrm{tr}(t_{\sigma}\cdot(\iota(v_{1}\otimes\lambda_{1})\otimes\iota(v_{2}\otimes\lambda_{2})\otimes\iota(v_{3}\otimes\lambda_{3})))$
$=$ $\mathrm{tr}(t_{\sigma}\cdot\iota(v_{1}\otimes v_{2}\otimes v_{3}\otimes\lambda_{1}\otimes\lambda_{2}\otimes\lambda_{3}))$
$=$ $\mathrm{tr}(\iota(t_{\sigma}(v_{1}\otimes v_{2}\otimes v_{3})\otimes\lambda_{1}\otimes\lambda_{2}\otimes\lambda_{3}))$
$=$ $\mathrm{tr}(\iota(v_{\sigma(1)}\otimes v_{\sigma(2)}\otimes v_{\sigma(3)}\otimes\lambda_{1}\otimes\lambda_{2}\otimes\lambda_{3}))$
$=$ $\lambda_{1}(v_{\sigma(1)})\lambda_{2}(v_{\sigma(2)})\lambda_{3}(v_{\sigma(3)}).$
From here, the results follow for each of the three particular permutations $\sigma$ being considered. ∎
In the next section, in the proof of our main result (Proposition 18), we will need the following Lemma 10, which is about composing the $L_{t_{\sigma}}$ with forms that are simple tensors. In the proof of Lemma 10, we will need this simpler lemma about evaluating the $L_{t_{\sigma}}$ on simple tensors:
Lemma 9.
For any vector space $U$, for any $u_{1},u_{2},u_{3}$ in $U$, any $\zeta_{1},\zeta_{2},\zeta_{3}$ in $U^{*}$, and any permutation $\sigma$ in $S_{3}$, we have the following equality between elements of $\mathrm{L}(U^{\otimes 3})$:
$L_{t_{\sigma}}\left(\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})\right)=\bigotimes_{i=1,2,3}\iota(u_{\sigma(i)}\otimes\zeta_{i}).$
Proof.
We have
$L_{t_{\sigma}}\left(\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})\right)$ $=t_{\sigma}\cdot\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})$
$=\bigotimes_{i=1,2,3}\iota(t_{\sigma}(u_{i})\otimes\zeta_{i})$
$=\bigotimes_{i=1,2,3}\iota(u_{\sigma(i)}\otimes\zeta_{i}).\;\qed$
Lemma 10.
For any vector space $U$, for any $v_{1},v_{2},v_{3}$ in $U$, any $\lambda_{1},\lambda_{2},\lambda_{3}$ in $U^{*}$, and any permutation $\sigma$ in $S_{3}$, we have the following equality between elements of $\mathrm{L}(U^{\otimes 3})^{*}$:
$\left(\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{i})\right)\circ L_{t_{\sigma}}=\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{\sigma^{-1}(i)}).$
Proof.
By linearity, it is enough to verify that both sides agree when evaluated on a rank one tensor of the form $\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})$ for some $u_{i}$ in $U$ and $\zeta_{i}$ in $U^{*}$. We have:
$\left(\left(\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{i})\right)\circ L_{t_{\sigma}}\right)\left(\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})\right)$
$=\left(\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{i})\right)\left(\bigotimes_{i=1,2,3}\iota(u_{\sigma(i)}\otimes\zeta_{i})\right)$ by Lemma 9
$=\prod_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{i})(\iota(u_{\sigma(i)}\otimes\zeta_{i}))$
[…]
$=\prod_{i=1,2,3}\lambda_{\sigma^{-1}(i)}(u_{i})\zeta_{i}(v_{i})$ by commutativity in $k$
[…]
$=\left(\bigotimes_{i=1,2,3}\iota^{*}(v_{i}\otimes\lambda_{\sigma^{-1}(i)})\right)\left(\bigotimes_{i=1,2,3}\iota(u_{i}\otimes\zeta_{i})\right).$
4 Main results
Let us return to the 2-dimensional vector space $V$ over a field $k$ that we had fixed in the overview. Let $I$ denote the identity in $\mathrm{L}(V)$.
Definition 11.
Let $Q=\mathrm{L}(V)/kI$ denote the quotient vector space of $\mathrm{L}(V)$ by the scalar multiples of identity.
Note that $\dim Q=(\dim V)^{2}-1=3$. The dual $Q^{*}$ is identified with the subspace of $\mathrm{L}(V)^{*}$ consisting of those forms $\mu$ that satisfy $\mu(I)=0$.
Lemma 12.
The trilinear form $g$ on $\mathrm{L}(V)$ is antisymmetric and passes to the quotient $Q$, inducing a volume form on $Q$.
Proof.
The antisymmetry follows from the definition of $g$ in Equation (2) and the cyclic property of the trace, $\mathrm{tr}(a_{1}a_{2}a_{3})=\mathrm{tr}(a_{2}a_{3}a_{1})$. The claim about passing to the quotient is that for $a_{1},a_{2},a_{3}$ in $\mathrm{L}(V)$, if any of the $a_{i}$ is a scalar multiple of identity, then $g(a_{1},a_{2},a_{3})=0$. This is verified directly, for instance if $a_{3}=I$ then $g(a_{1},a_{2},a_{3})=\mathrm{tr}(a_{1}a_{2})-\mathrm{tr}(a_{2}a_{1})=0$. Finally, as $\dim Q=3$, antisymmetric 3-forms on $Q$ are volume forms on $Q$. ∎
Lemma 13.
For any basis $(\mu_{1},\mu_{2},\mu_{3})$ of $Q^{*}$, letting $\varepsilon(\sigma)$ denote the signature of a permutation $\sigma$, there exists a scalar $\alpha$ such that
$g=\alpha\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\bigotimes_{i=1,2,3}\mu_{\sigma(i)}.$ (10)
Proof.
As the space of volume forms on $Q$ is one-dimensional, and by Lemma 12 we already know that $g$ is a volume form on $Q$, it is enough to check that the right-hand side is antisymmetric. That is true by construction, that expression being known as an antisymmetrized tensor product. ∎
Remark 14.
The constant $\alpha$ in Lemma 13 can be computed by picking any $c_{1},c_{2},c_{3}$ in $\mathrm{L}(V)$ such that $g(c_{1},c_{2},c_{3})=1$ and using Equation (10) as a definition of $\alpha^{-1}$:
$\alpha^{-1}=\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\prod_{i=1,2,3}\mu_{\sigma(i)}(c_{i}).$ (11)
Lemma 15.
The following equalities hold between trilinear forms on $\mathrm{L}(V)$:
$g$ $=$ $t_{(123)}^{*}-t_{(321)}^{*},$ (12)
$h$ $=$ $t_{\mathrm{id}}^{*}-t_{(123)}^{*}.$ (13)
Proof.
This follows readily from Lemma 8 and the definitions of $g$ and $h$ in Equations (2, 3). ∎
Lemma 16.
The following equality holds between forms in $\mathrm{L}(V^{\otimes 3})^{*}$:
$h=g\circ L_{t_{(321)}}.$
Proof.
We have
$h$ $=(t_{\mathrm{id}}-t_{(123)})^{*}$ by Equation (13)
$=((t_{(123)}-t_{(321)})\cdot t_{(321)})^{*}$
$=(t_{(123)}-t_{(321)})^{*}\circ L_{t_{(321)}}$ by Lemma 6
$=g\circ L_{t_{(321)}}$ by Equation (12).∎
While Lemma 13 allowed arbitrary linear forms $\mu_{i}$, Proposition 18 will need to restrict to rank one forms, meaning the $\iota^{*}(v\otimes\lambda)$ for $v$ in $V$ and $\lambda$ in $V^{*}$. The necessity of that restriction is discussed in Remark 19.
Lemma 17.
For $i=1,2,3$, let $v_{i}$ be a nonzero vector in $V$, let $\lambda_{i}$ be a nonzero linear form on $V$ such that $\lambda_{i}(v_{i})=0$, and let $\mu_{i}=\iota^{*}(v_{i}\otimes\lambda_{i})$. The following conditions are equivalent:
  1. 1.
    The vectors $v_{1},v_{2},v_{3}$ are pairwise noncolinear: $i\neq j\Rightarrow v_{i}\not\in\mathrm{span}(v_{j})$.
  2. 2.
    The forms $\lambda_{1},\lambda_{2},\lambda_{3}$ are pairwise noncolinear: $i\neq j\Rightarrow\lambda_{i}\not\in\mathrm{span}(\lambda_{j})$.
  3. 3.
    The forms $\mu_{1},\mu_{2},\mu_{3}$ are linearly independent.
  4. 4.
    The family $(\mu_{1},\mu_{2},\mu_{3})$ is a basis of $Q^{*}$.
Proof.
[…]
$\mathrm{span}(v_{i})=\ker(\lambda_{i}),$
since $\lambda_{i}(v_{i})=0$ means that $\mathrm{span}(v_{i})\subset\ker(\lambda_{i})$, and as $\dim V=2$, we have $\dim\ker(\lambda_{i})=1$ hence the inclusion is an equality. This means in particular that for each i, j,
[…]
Proposition 18.
For any vectors $v_{1},v_{2},v_{3}$ in $V$ and linear forms $\lambda_{1},\lambda_{2},\lambda_{3}$ on $V$ satisfying the equivalent conditions of Lemma 17, we have:
$h=\frac{-1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\bigotimes_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma(123)(i)}).$ (14)
Proof.
Let us first explain why the denominator $\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})$ is nonzero. Because of condition 1 in Lemma 17, whenever $i\neq j$, the vector $v_{j}$ cannot belong to the one-dimensional space $\ker(\lambda_{i})=\mathrm{span}(v_{i})$, so $\lambda_{i}(v_{j})\neq 0$, so $\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})\neq 0$.
Let us now prove Equation (14) up to a scalar factor $\alpha$. Let $\mu_{i}=\iota^{*}(v_{i}\otimes\lambda_{i})$. Lemma 17 says that the $\mu_{i}$ form a basis of $Q^{*}$, so we can apply Lemma 13 with that basis to obtain
$g=\alpha\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\bigotimes_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma(i)})$
for some scalar $\alpha$. Lemma 16 transforms that into
$h=\alpha\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\left(\bigotimes_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma(i)})\right)\circ L_{t_{(321)}},$
which Lemma 10 transforms into
$h=\alpha\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\bigotimes_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma(123)(i)}).$ (15)
There only remains to evaluate the scalar $\alpha$. Let $a_{i}=\iota(v_{i}\otimes\lambda_{i})$. Notice that $\mathrm{tr}(a_{i})=\lambda_{i}(v_{i})=0$, so
$h(a_{1},a_{2},a_{3})=-tr(a_{1}a_{2}a_{3})=-\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1}).$ (16)
On the other hand, evaluating Equation (15) and simplifying that using Equation (6) yields
$h(a_{1},a_{2},a_{3})=\alpha\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\prod_{i=1,2,3}\lambda_{\sigma(123)(i)}(v_{i})\lambda_{i}(v_{\sigma(i)}).$ (17)
Since $\lambda_{i}(v_{i})=0$, the product in Equation (17) vanishes whenever $\sigma$ has a fixed point or $\sigma(123)$ has a fixed point. Thus the only $\sigma$ contributing to the sum is $\sigma=(123)$. Thus, Equation (17) simplifies to
$h(a_{1},a_{2},a_{3})=\alpha\prod_{i=1,2,3}\lambda_{(321)(i)}(v_{i})\lambda_{i}(v_{(123)(i)}).$ (18)
further simplifying as
$h(a_{1},a_{2},a_{3})=\alpha(\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1}))^{2}.$
Combining that with Equation (16) yields
$\alpha=\frac{-1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\cdot\;\qed$
Remark 19.
Two tensors $p,q$ in $\mathrm{L}(V)^{*\otimes 3}$ related to each other in the same way as $g$ and $h$ are related by Lemma 16, namely $q=p\circ L_{t_{(321)}}$, may still fail to have the same tensor rank if their tensor decompositions involve linear terms in $\mathrm{L}(V)^{*}$ that are not of rank one.
Proof.
Consider the counterexample of $p=t_{(123)}^{*}$ and $q=t_{\mathrm{id}}^{*}$. The same argument as in the proof of Lemma 16 yields $q=p\circ L_{t_{(321)}}$. As noted in Lemma 8, for $a_{1},a_{2},a_{3}$ in $\mathrm{L}(V)$, we have $p(a_{1},a_{2},a_{3})=\mathrm{tr}(a_{1}a_{2}a_{3})$ and $q(a_{1}a_{2}a_{3})=\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})$. Thus, as tensors in $\mathrm{L}(V)^{*\otimes 3}$, $q$ has rank one but $p$ does not. ∎
To elaborate on the previous remark, the linear form $a\mapsto\mathrm{tr}(a)$ does not have rank one, so even though $q$ has rank one as a tensor of order 3 in $\mathrm{L}(V)^{*\otimes 3}$, it does not have rank one as a tensor of order 6 in $(V\otimes V^{*})^{\otimes 3}$, and our tool for transporting tensor decompositions, Lemma 10, applies to tensors of order 6 in $(V\otimes V^{*})^{\otimes 3}$.
5 Strassen algorithms
Proposition 18 is already a form of Strassen’s algorithm, but that may be obscured by the tensor formalism, so let us derive a few more concrete statements as corollaries.
Corollary 20.
For any vectors $v_{1},v_{2},v_{3}$ in $V$ and linear forms $\lambda_{1},\lambda_{2},\lambda_{3}$ on $V$ satisfying the equivalent conditions of Lemma 17, for all $a_{1},a_{2},a_{3}$ in $\mathrm{L}(V)$,
$\mathrm{tr}(a_{1}a_{2}a_{3})=\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})\\ +\frac{1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\prod_{i=1,2,3}\lambda_{\sigma(123)(i)}(a_{i}(v_{\sigma(i)})).$
Proof.
Evaluating Equation (14) at any $a_{1},a_{2},a_{3}$ in $\mathrm{L}(V)$ gives:
$\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})\mathrm{tr}(a_{3})-\mathrm{tr}(a_{1}a_{2}a_{3})=\\ \frac{-1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\prod_{i=1,2,3}\iota^{*}(v_{\sigma(i)}\otimes\lambda_{\sigma(123)(i)})(a_{i})$
and the result follows by Definition 1. ∎
Corollary 21.
For any vectors $v_{1},v_{2},v_{3}$ in $V$ and linear forms $\lambda_{1},\lambda_{2},\lambda_{3}$ on $V$ satisfying the equivalent conditions of Lemma 17, for all $a_{1},a_{2}$ in $\mathrm{L}(V)$,
$a_{1}a_{2}=\mathrm{tr}(a_{1})\mathrm{tr}(a_{2})I\\ +\frac{1}{\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})}\sum_{\sigma\in S_{3}}\varepsilon(\sigma)\mathrm{tr}(a_{1}c_{\sigma(1),\sigma(2)})\mathrm{tr}(a_{2}c_{\sigma(2),\sigma(3)})c_{\sigma(3),\sigma(1)}$ (19)
where $c_{i,j}$ in $\mathrm{L}(V)$ is defined by $c_{i,j}(u)=\lambda_{j}(u)v_{i}$ for all $u$ in $V$.
Proof.
Let $x$ denote the right-hand side of Equation (19). The claim is that $a_{1}a_{2}=x$. That is equivalent to the claim that $\mathrm{tr}(a_{1}a_{2}a_{3})=\mathrm{tr}(xa_{3})$ for all $a_{3}$ in $\mathrm{L}(V)$. That claim is directly verified by comparing the expression of $\mathrm{tr}(a_{1}a_{2}a_{3})$ given by Corollary 20 to the expression of $\mathrm{tr}(xa_{3})$ expanded by using the definition of $x$, noting that $c_{i,j}=\iota(v_{i}\otimes\lambda_{j})$. ∎
Corollary 22.
The original Strassen algorithm is obtained by applying Corollary 21 to the vector space $V=k^{2}$, with the following choices: $v_{1}=\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$, $\lambda_{1}=\left(\begin{smallmatrix}0&1\end{smallmatrix}\right)$, $v_{2}=\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)$, $\lambda_{2}=\left(\begin{smallmatrix}1&0\end{smallmatrix}\right)$, $v_{3}=\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$, $\lambda_{3}=\left(\begin{smallmatrix}1&-1\end{smallmatrix}\right)$.
Proof.
Applying Corollary 21, expanding the sum over all 6 permutations, and noticing that $\lambda_{1}(v_{2})\lambda_{2}(v_{3})\lambda_{3}(v_{1})=1$, we obtain the following matrix multiplication algorithm: for any two $2\times 2$ matrices $a,b$,
$ab$ $=\mathrm{tr}(a)\mathrm{tr}(b)I$
$+\mathrm{tr}(ac_{1,2})\mathrm{tr}(bc_{2,3})c_{3,1}$
$+\mathrm{tr}(ac_{2,3})\mathrm{tr}(bc_{3,1})c_{1,2}$
$+\mathrm{tr}(ac_{3,1})\mathrm{tr}(bc_{1,2})c_{2,3}$
$-\mathrm{tr}(ac_{2,1})\mathrm{tr}(bc_{1,3})c_{3,2}$
$-\mathrm{tr}(ac_{1,3})\mathrm{tr}(bc_{3,2})c_{2,1}$
$-\mathrm{tr}(ac_{3,2})\mathrm{tr}(bc_{2,1})c_{1,3}$
where the $c_{i,j}=\iota(v_{i}\otimes\lambda_{j})=v_{i}\lambda_{j}$ are:
$c_{1,2}=v_{1}\lambda_{2}=\left(\begin{matrix}1&0\\ 0&0\end{matrix}\right),$ $c_{1,3}=v_{1}\lambda_{3}=\left(\begin{matrix}1&-1\\ 0&0\end{matrix}\right),$
$c_{2,3}=v_{2}\lambda_{3}=\left(\begin{matrix}0&0\\ 1&-1\end{matrix}\right),$ $c_{2,1}=v_{2}\lambda_{1}=\left(\begin{matrix}0&0\\ 0&1\end{matrix}\right),$
$c_{3,1}=v_{3}\lambda_{1}=\left(\begin{matrix}0&1\\ 0&1\end{matrix}\right),$ $c_{3,2}=v_{3}\lambda_{2}=\left(\begin{matrix}1&0\\ 1&0\end{matrix}\right).$
Let $a^{i,j}$ and $b^{i,j}$ denote the matrix coefficients, using superscript notation to distinguish that from the subscripts used to index the $c_{i,j}$ matrices. Let $e_{i,j}$ be the elementary matrix with a 1 at position $(i,j)$ and zeros elsewhere. Using the above table of $c_{i,j}$ matrices, the above equation expands to
$ab$ $=(a^{1,1}+a^{2,2})(b^{1,1}+b^{2,2})(e_{1,1}+e_{2,2})$
$+a^{1,1}(b^{1,2}-b^{2,2})(e_{1,2}+e_{2,2})$
$+(a^{1,2}-a^{2,2})(b^{2,1}+b^{2,2})e_{1,1}$
$+(a^{2,1}+a^{2,2})b^{1,1}(e_{2,1}-e_{2,2})$
$-a^{2,2}(b^{1,1}-b^{2,1})(e_{1,1}+e_{2,1})$
$-(a^{1,1}-a^{2,1})(b^{1,1}+b^{1,2})e_{2,2}$
$-(a^{1,1}+a^{1,2})b^{2,2}(e_{1,1}-e_{1,2}).$
These bilinear forms in the $a^{i,j}$ and $b^{i,j}$ are exactly the terms I, II, III, IV, V, VI, VII introduced in the original Strassen article [1]:
$ab$ $=\mathrm{I}\cdot(e_{1,1}+e_{2,2})$
$+\mathrm{III}\cdot(e_{1,2}+e_{2,2})$
$+\mathrm{VII}\cdot e_{1,1}$
$+\mathrm{II}\cdot(e_{2,1}-e_{2,2})$
$+\mathrm{IV}\cdot(e_{1,1}+e_{2,1})$
$+\mathrm{VI}\cdot e_{2,2}$
$+\mathrm{V}\cdot(e_{1,2}-e_{1,1}).$
Thus the coefficients of the product matrix $ab$ are:
$(ab)^{1,1}$ $=\mathrm{I}+\mathrm{IV}-\mathrm{V}+\mathrm{VII}$
$(ab)^{1,2}$ $=\mathrm{III}+\mathrm{V}$
$(ab)^{2,1}$ $=\mathrm{II}+\mathrm{IV}$
$(ab)^{2,2}$ $=\mathrm{I}-\mathrm{II}+\mathrm{III}+\mathrm{VI}$
exactly as originally stated by Strassen [1]. ∎